OpenStudy (anonymous):

400 feet in 52.4 seconds a. 76 feet per second c. 7.7 feet per second b. 0.76 foot per second d. 7.6 feet per second how do i solve this?

OpenStudy (danjs):

\[\frac{ 400~feet }{ 52.4~second } = \frac{ x ~ft }{ 1~second }\]

OpenStudy (anonymous):

cross multiply?

OpenStudy (danjs):

yes, you can do that

OpenStudy (danjs):

all you really have to do is multiply both sides by 1 second

OpenStudy (danjs):

divide 400 by 52.4

OpenStudy (anonymous):

D

OpenStudy (danjs):

\[\frac{ 400~feet }{ 52.4~second }=\frac{ (400/52.4)~ft }{ 1 ~s }\]